# Geometry-Aware Spiking Graph Neural Network

### v1

@@ Introduction
Graphs are ubiquitous non-Euclidean structures used to model complex relationships in real-world systems. Graph Neural Networks (GNNs) (Hamilton, Ying, and Leskovec 2017; Ju et al. 2024; Kipf and Welling 2017a), built upon floating-point Artificial Neural Networks (ANNs), have achieved impressive success in learning expressive graph representations. However, their performance often comes at the cost of high computational and energy demands, particularly when scaling to large and complex graphs (Zhu et al. 2022; Yin et al. 2024c). In parallel, Spiking Neural Networks (SNNs), inspired by the spike-based communication mechanism of biological neurons, offer event-driven and temporally sparse computation, demonstrating notable advantages in energy efficiency (Maass 1997; Yin et al. 2024b). Combining the structural modeling capability of GNNs with the energy efficiency of SNNs, spiking GNNs are rapidly gaining attention as a promising new paradigm in graph learning.
Despite recent advances in spiking GNNs, their representational expressiveness and adaptability remain fundamentally constrained by two key limitations. Prior efforts have primarily implemented spiking GNNs within Euclidean space (Sarkar 2011), inheriting the design of conventional artificial neural networks. Although this design is computationally convenient, it struggles to capture the intricate geometric structures present in many real-world graphs. For instance, hierarchical social networks (Wang and Jiang 2022) and molecular graphs with ring-like dependencies (Yin et al. 2023d; Yang et al. 2019) naturally exhibit non-Euclidean characteristics, which cannot be accurately embedded into flat space without significant distortion. While hyperbolic and spherical manifolds have demonstrated greater suitability for preserving such structural properties (Chami et al. 2019; Yin et al. 2023b), their integration into spiking GNNs has received little attention. Furthermore, most existing approaches assume a fixed geometric prior for the entire dataset, whether Euclidean or based on a specific manifold (Yin et al. 2023c; Bachmann, Bécigneul, and Ganea 2020; Pang et al. 2023). This global assumption overlooks the structural diversity across individual graphs and limits the model’s ability to adapt to varying local geometries. Although a few recent studies propose using mixed-curvature product spaces to improve representational flexibility (Wang et al. 2023; Yin et al. 2023a), they adopt a shared geometric configuration across all samples. As a result, these methods are unable to adjust their geometric inductive biases in an instance-aware manner. These limitations highlight the need for a principled framework that can dynamically select the most appropriate geometry for each input graph, enabling spiking GNNs to more effectively align with the intrinsic structure of the data.
@@
Spiking Neural Networks (SNNs) (Yin et al. 2022b; Ai et al. 2023; Yin and Luo 2022) draw inspiration from the firing patterns of biological neurons and achieve energy-efficient computation by transmitting information through sparse, binary spike signals in an event-driven manner. They offer significant energy advantages over traditional artificial neural networks, making them attractive for neuromorphic hardware and low-power applications (Yin et al. 2022a; Tavanaei et al. 2019). The integration of SNNs with graph structures has given rise to spiking GNNs, which adapt the spike-based computation paradigm to graph data (Shou et al. 2023; Zhu et al. 2022). Early designs focus on replacing continuous activations with spiking neurons in standard GNN architectures such as convolutional or attention-based models (Tang et al. 2024; Zhu et al. 2022; Yin et al. 2024a). While promising, these approaches often rely on surrogate gradients and recurrent training schemes like Backpropagation Through Time (BPTT) (Zhang et al. 2022), which incur high computational latency. More recently, researchers have sought to enhance both efficiency and representational power by introducing geometric awareness. The Manifold-valued Spiking GNN (Sun et al. 2024a) introduces a novel spiking layer operating in Riemannian manifolds rather than Euclidean space, leveraging geodesic dynamics and diffeomorphisms to bridge spike trains and manifold representations. Crucially, it bypasses the BPTT bottleneck via a training paradigm called Differentiation via Manifold, enabling recurrence-free gradient propagation. However, current methods typically restrict representation learning to a single manifold, overlooking the diverse geometric structures in real-world graphs and thus limiting their expressivity.
@@
GNNs in Euclidean and Riemannian Spaces. Graph Neural Networks (GNNs) have become foundational tools for learning representations from graph-structured data (Zhang et al. 2023; Wu et al. 2019; Meng et al. 2024). Traditional GNNs typically operate in Euclidean space, which assumes flat geometry and offers computational simplicity. However, this assumption often leads to distortion when modeling real-world graphs with hierarchical, cyclical, or heterogeneous structures. To address these challenges, researchers have explored non-Euclidean spaces. Hyperbolic spaces, with negative curvature, are well-suited for hierarchical structures due to their exponential capacity, while spherical spaces, with positive curvature, better capture cyclic or angular patterns (Coors, Condurache, and Geiger 2018). These insights have motivated the development of Riemannian GNNs (Ding et al. 2024; Zhang et al. 2025), which generalize neural operations to curved manifolds such as hyperbolic, spherical, and their product spaces. Recent methods further propose mixed-curvature manifolds to better model structural diversity. However, most existing approaches adopt a fixed geometric space for all data, ignoring instance-level geometric variability. Our method addresses this limitation by dynamically selecting and combining multiple Riemannian spaces based on the structural properties of each input, resulting in more flexible and expressive graph representations.

### v2

@@ Introduction
Graphs are ubiquitous non-Euclidean structures used to model complex relationships in real-world systems. Graph Neural Networks (GNNs) (Hamilton, Ying, and Leskovec 2017; Veličković et al. 2018; Kipf and Welling 2017a), built upon floating-point Artificial Neural Networks (ANNs), have achieved impressive success in learning expressive graph representations. However, their performance often comes at the cost of high computational and energy demands, particularly when scaling to large and complex graphs (Zhu et al. 2022; Yin et al. 2024). In parallel, Spiking Neural Networks (SNNs), inspired by the spike-based communication mechanism of biological neurons, offer event-driven and temporally sparse computation, demonstrating notable advantages in energy efficiency (Maass 1997; Brette et al. 2007). Combining the structural modeling capability of GNNs with the energy efficiency of SNNs, spiking GNNs are rapidly gaining attention as a promising new paradigm in graph learning.
Despite recent advances in spiking GNNs, their representational expressiveness and adaptability remain fundamentally constrained by two key limitations. Prior efforts have primarily implemented spiking GNNs within Euclidean space (Sarkar 2011), inheriting the design of conventional artificial neural networks. Although this design is computationally convenient, it struggles to capture the intricate geometric structures present in many real-world graphs. For instance, hierarchical social networks (Wang and Jiang 2022) and molecular graphs with ring-like dependencies (Gilmer et al. 2017; Yang et al. 2019) naturally exhibit non-Euclidean characteristics, which cannot be accurately embedded into flat space without significant distortion. While hyperbolic and spherical manifolds have demonstrated greater suitability for preserving such structural properties (Chami et al. 2019; Xiong et al. 2022), their integration into spiking GNNs has received little attention. Furthermore, most existing approaches assume a fixed geometric prior for the entire dataset, whether Euclidean or based on a specific manifold (Chen et al. 2021; Bachmann, Bécigneul, and Ganea 2020; Xiong et al. 2022). This global assumption overlooks the structural diversity across individual graphs and limits the model’s ability to adapt to varying local geometries. Although a few recent studies propose using mixed-curvature product spaces to improve representational flexibility (Wang et al. 2023; Zhang and Jiang 2023), they adopt a shared geometric configuration across all samples. As a result, these methods are unable to adjust their geometric inductive biases in an instance-aware manner. These limitations highlight the need for a principled framework that can dynamically select the most appropriate geometry for each input graph, enabling spiking GNNs to more effectively align with the intrinsic structure of the data.
@@
Spiking Neural Networks (SNNs) (Maass 1997; Brette et al. 2007; Cao, Chen, and Khosla 2015) draw inspiration from the firing patterns of biological neurons and achieve energy-efficient computation by transmitting information through sparse, binary spike signals in an event-driven manner. They offer significant energy advantages over traditional artificial neural networks, making them attractive for neuromorphic hardware and low-power applications (Davies et al. 2018; Tavanaei et al. 2019). The integration of SNNs with graph structures has given rise to spiking GNNs, which adapt the spike-based computation paradigm to graph data (Xu et al. 2021; Zhu et al. 2022). Early designs focus on replacing continuous activations with spiking neurons in standard GNN architectures such as convolutional or attention-based models (Li et al. 2023a; Zhu et al. 2022; Sun et al. 2024b). While promising, these approaches often rely on surrogate gradients and recurrent training schemes like Backpropagation Through Time (BPTT) (Huh and Sejnowski 2018), which incur high computational latency. More recently, researchers have sought to enhance both efficiency and representational power by introducing geometric awareness. The Manifold-valued Spiking GNN (Sun et al. 2024a) introduces a novel spiking layer operating in Riemannian manifolds rather than Euclidean space, leveraging geodesic dynamics and diffeomorphisms to bridge spike trains and manifold representations. Crucially, it bypasses the BPTT bottleneck via a training paradigm called Differentiation via Manifold, enabling recurrence-free gradient propagation. However, current methods typically restrict representation learning to a single manifold, overlooking the diverse geometric structures in real-world graphs and thus limiting their expressivity.
@@
GNNs in Euclidean and Riemannian Spaces. Graph Neural Networks (GNNs) have become foundational tools for learning representations from graph-structured data (Kipf and Welling 2017b; Wu et al. 2019; Veličković et al. 2018). Traditional GNNs typically operate in Euclidean space, which assumes flat geometry and offers computational simplicity. However, this assumption often leads to distortion when modeling real-world graphs with hierarchical, cyclical, or heterogeneous structures. To address these challenges, researchers have explored non-Euclidean spaces. Hyperbolic spaces, with negative curvature, are well-suited for hierarchical structures due to their exponential capacity, while spherical spaces, with positive curvature, better capture cyclic or angular patterns (Coors, Condurache, and Geiger 2018). These insights have motivated the development of Riemannian GNNs (Chami et al. 2019; Liu, Nickel, and Kiela 2019), which generalize neural operations to curved manifolds such as hyperbolic, spherical, and their product spaces. Recent methods further propose mixed-curvature manifolds to better model structural diversity. However, most existing approaches adopt a fixed geometric space for all data, ignoring instance-level geometric variability. Our method addresses this limitation by dynamically selecting and combining multiple Riemannian spaces based on the structural properties of each input, resulting in more flexible and expressive graph representations.